# Authors Should Annotate

### v1

@@ 5.3 Internal Consistency and Predictivity
Alignment with author labels holds no meaning if the author labels themselves do not carry any useful signal. We conduct an experiment to see how internally consistent and predictive different annotation types are, including author labeling. In this setup, we prompt an LLM to predict what label the user would assign a question with a similar prompt as the LLM annotator’s (Appendix C.2); however, we also provide as context five other questions asked to the same user along with labels from five different sources: the author label, the three 3rd-party annotation methods, and random labels. Our intuition is that the predictor will perform better with the additional information given from other questions, and that the higher quality the label source, the better the performance.
@@ Results
We also explored whether it mattered if the task examples occurred in the same conversation in Figure 6. We find that some methods benefit from a higher number of in-conversation task answers while others do not, but also note that this differs by model as well (See Appendix C.3. However, across all models, author labeling benefits the most from additional in-conversation examples, as intuitively additional context from the examples is likely most relevant in the same conversation and timespan of the actual question and author labeling contains the highest fidelity of annotator answer.
@@ 5.4 Cost Analysis of Annotation Methods
We conduct a simple cost and time analysis comparing author labeling to the two human annotation methods, Mechanical Turk crowdsourcing and expert annotation. We rewarded chatbot users with \$0.08 per task response, with each task taking less than ten seconds to complete (five seconds of forced response delay plus five seconds to answer). This yields a rate of \$28.80 per hour for author labeling tasks. For experts, while they graciously volunteered for free, we consider a rate of \$20 per hour to be reasonable. For Mechanical Turk, we offered a fixed rate of \$5 for a 30-question survey, which we initially estimated to take 30 minutes. In actuality, it took Mechanical Turk annotators an average of 9.1 minutes to complete the survey and expert annotators 17.0 minutes [^5] The 9.1 minute survey completion for MTurk likely indicates these annotators did not entirely read through each conversation, as assuming a fast reading rate of 250 WPM, this would mean they read 50 words per conversation, well below the 125-140 word average.. We present the cost per datum and time per datum in Table 4, which shows author labeling as significantly cheaper and faster than the other annotation methods. We note that our calculations do not factor in that 3rd-party annotation methods typically require multiple annotators per datum; these rates would double or triple when considering additional annotators.
@@ 7 Ethics
In order to protect users’ privacy, we will not release any of the personal conversations used in our analysis. We also redact names, locations, and other PII when presenting it to our annotators and omit any conversations dealing with sensitive topics such as personal health or finances. Users fully consent to their data being analyzed for business and scientific purposes when signing up to echo LLM and are fully informed on how they will be compensated for filing out label questions. Filling out label questions is always completely optional and their decision to fill out or not fill out a task does not affect their future experience on echo in any way.
@@ 8 Limitations
We previously discussed the theoretical limitations of author labeling in §2.3, but now discuss additional experimental limitations. Because of the real-world nature of our experiments, we could not control for all factors that occurred in the chatbot. The userbase itself was growing throughout the four-week trial period, and some spammy users were banned or excluded from analysis after the fact. Additionally, our head-to-head analysis of 3rd-party annotation to author labeling dealt with egocentric sentiment analysis. We did not analyze objective annotations or non-egocentric subjective tasks, both of which are tasks where being the author of the document might not be as important to annotation. In these scenarios, the qualifications of the author matter, which were factors that were outside the scope of this work. We also used LLMs to generate the questions in order to be as contextually relevant as possible, but the phrasing of questions can greatly influence self-reports schwarz_1999. Additionally, although we introduce the general architectural framework of Echo, we only test its efficacy for the downstream task of product recommendation.

### v2

@@ 5.3 Internal Consistency and Predictivity
Alignment with author labels holds no meaning if the author labels themselves do not carry any useful signal. We conduct an experiment to see how internally consistent and predictive different annotation types are, including author labeling. Our goal is to use a user’s prior answers to predict a future answer. In this setup, we prompt an LLM to predict what label the user would assign a question (Appendix C.2); however, we also provide as context five other questions previously asked to the same user along with their labels from five different sources: the author label, the three 3rd-party annotation methods, and random labels. Our intuition is that the predictor will perform better with the additional information given from other questions, and that the higher quality the label source, the better the performance.
@@ Results
We also explored whether it mattered if the task examples occurred in the same conversation in Figure 6. We find that some methods benefit from a higher number of in-conversation task answers while others do not, but also note that this differs by model as well (See Appendix C.3). However, across all models, author labeling benefits the most from additional in-conversation examples, as intuitively additional context from the examples is likely most relevant in the same conversation and timespan of the actual question and author labeling contains the highest fidelity of annotator answer.
@@ 5.4 Cost Analysis of Annotation Methods
We conduct a cost and time analysis comparing author labeling to the two human annotation methods, Mechanical Turk crowdsourcing and expert annotation. We rewarded chatbot users with \$0.08 per task response, with each task taking less than ten seconds to complete (five seconds of forced response delay plus five seconds to answer). This yields a rate of \$28.80 per hour for author labeling tasks. For experts, while they graciously volunteered for free, we consider a rate of \$20 per hour to be reasonable. For Mechanical Turk, we offered a fixed rate of \$5 for a 30-question survey, which we initially estimated to take 30 minutes. In actuality, it took Mechanical Turk annotators an average of 9.1 minutes to complete the survey and expert annotators 17.0 minutes [^6] The 9.1 minute survey completion for MTurk likely indicates these annotators did not entirely read through each conversation, as assuming a fast reading rate of 250 WPM, this would mean they read 50 words per conversation, well below the 125-140 word average.. We present the cost per datum and time per datum in Table 4, which shows author labeling as significantly cheaper and faster than the other annotation methods. We note our calculations assume a single annotation per datum; these rates would double or triple when adopting conventional multiple-annotator coverage.
@@ 7 Ethics
In order to protect users’ privacy, we will not release any of the personal conversations used in our analysis besides the handful of non-identifiable examples used throughout this paper. We also redacted names, locations, and other PII when presenting it to our annotators and omitted any conversations dealing with sensitive topics such as personal health, religion, or finances. Users fully consented to their data being analyzed for business and scientific purposes when signing up to the chatbot and were fully informed on how they will be compensated for filing out label questions. Filling out label questions is always completely optional and a user’s decision to fill out or not fill out a task does not affect their future experience on the platform in any way.
@@ 8 Limitations
We previously discussed the theoretical limitations of author labeling in §2.3, but now discuss additional experimental limitations. Because of the real-world nature of our experiments, we could not control for all factors that occurred in the chatbot. The userbase itself was growing throughout the four-week trial period, and some spammy users were banned or excluded from analysis after the fact. Additionally, our head-to-head analysis of 3rd-party annotation to author labeling dealt with egocentric sentiment analysis. We did not analyze objective annotations or non-egocentric subjective tasks, both of which are tasks where being the author of the document might not be as important to annotation. In these scenarios, the qualifications of the author matter, which were factors that were outside the scope of this work. We also used LLMs to generate the questions in order to be as contextually relevant as possible, but the phrasing of questions can greatly influence self-reports schwarz_1999; abdurahman2024perils. Additionally, although we introduce the general architectural framework of Echo, we only test its efficacy for the task of product recommendation.